Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 129,000 $ 132,000
Accounts receivable 340,000 479,000
Inventory 577,000 476,000
Plant and equipment, net 794,000 793,000
Investment in Buisson, S.A. 401,000 427,000
Land (undeveloped) 255,000 249,000
Total assets $ 2,496,000 $ 2,556,000
Liabilities and Stockholders' Equity
Accounts payable $ 389,000 $ 344,000
Long-term debt 971,000 971,000
Stockholders' equity 1,136,000 1,241,000
Total liabilities and stockholders' equity $ 2,496,000 $ 2,556,000


Joel de Paris, Inc.
Income Statement
Sales $ 3,906,000
Operating expenses 3,359,160
Net operating income 546,840
Interest and taxes:
Interest expense $ 115,000
Tax expense 194,000 309,000
Net income $ 237,840


The company paid dividends of $132,840 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.)

3. What was the company’s residual income last year?

11. 14.00% Average operating assets Margin Turnover ROI 3. Residual income
Answer #1

Average operating assets

= [(2,496,000-255,000-401,000)+(2,556,000-249,000-427,000)]/2

= 1,860,000

Margin = Net operating income/Sales

= 546,840/3,906,000

= 14%

Turnover = Sales/Average operating assets

= 3,906,000/1,860,000

= 2.1

ROI = Margin * turnover

= 14*2.1

= 29.4%

Residual income = Net operating income - required return

= 546,840 - (1,860,000*15%)

= 267,840